OpenStudy (anonymous):

Find the general solution to the following differential equations f′′ − 4f′ + 4f = −4 + 12x.

myininaya (myininaya):

Don't forget to also solve f''-4f'+4f=0 :)

OpenStudy (anonymous):

The solution is \[ f(x) =A e^{2 x}+B x e^{2 x} +3 x+2 \] To get there first you solve the homegeneous equation \[ r^2 -4r + 4=(r-2)^2 \] r = 2 is adouble root. So the homegeneous equation has the general solution \[ f(x) =A e^{2 x}+B x e^{2 x} \] Now find a particualr solution of the general equation and add it to the other one to obatain \[ f(x) =A e^{2 x}+B x e^{2 x} +3 x+2 \]

OpenStudy (anonymous):

is there a way to do this through linear algebra?

OpenStudy (anonymous):

Yes, read http://tutorial.math.lamar.edu/Classes/DE/SystemsDE.aspx and try do it yourself.

OpenStudy (anonymous):

Let yp= c x + d be a particular solution we need to determine c and d. yp'= c and yp''= 0 replace in f′′ − 4f′ + 4f = −4 + 12x. to get 0 - 4c + 4 c x + 4 d = -4 + 12 x -4 c + 4d + 4 c x = -4 + 4 x 4c =12 ( coef. of x)=> c =3 -4c + 4 d = -4 => 12 + 4 d =-4 => 4 d= 8=> d = 2 so yp= 3 x + 2